mary who is 16 years old,is four times as old as her brother.how old will mary be when she is twice as old as her brother.?
Answer by stanbon(75887)

mary who is 16 years old,is four times as old as her brother.how old will mary be when she is twice as old as her brother.?
---
m = 16 yrs
b = 4 yrs
----
Equation:
Solve: 16+x = 2(4+x)
16+x = 8+2x
x = 8
---
Mary will be 16+8 = 24 yrs.
Brother will be 4+8 = 12 yrs.
============================
Cheers,
Stan H.